OpenStudy (anonymous):

Find the exact value of tan (arcsin (two fifths)). For full credit, explain your reasoning

OpenStudy (anonymous):



OpenStudy (anonymous):

Calculate the length of the third side.\[\sqrt{5^2-2^2}=\sqrt{21} \]The tangent is\[\frac{2}{\sqrt{21}} \]